### v2

# BigTokDetect: A Clinically-Informed Vision–Language Modeling Framework for Detecting Pro-Bigorexia Videos on TikTok

Abstract
Social media platforms increasingly struggle to detect harmful content that promotes muscle dysmorphic behaviors, particularly pro-bigorexia content that disproportionately affects adolescent males. Unlike traditional eating disorder detection focused on the "thin ideal," pro-bigorexia material masquerades as legitimate fitness content through complex multimodal combinations of visual displays, coded language, and motivational messaging that evade text-based detection systems. We address this challenge by developing BigTokDetect, a clinically-informed detection framework for identifying pro-bigorexia content on TikTok. We introduce BigTok, the first expert-annotated multimodal dataset of over 2,200 TikTok videos labeled by clinical psychologists and psychiatrists across five primary categories spanning body image, nutrition, exercise, supplements, and masculinity. Through comprehensive evaluation of state-of-the-art vision language models, we achieve 82.9% accuracy on primary category classification and 69.0% on subcategory detection via domain-specific finetuning. Our ablation studies demonstrate that multimodal fusion improves performance by 5-10% over text-only approaches, with video features providing the most discriminative signals. These findings establish new benchmarks for multimodal harmful content detection and provide both the computational tools and methodological framework needed for scalable content moderation in specialized mental health domains.
BigTokDetect: A Clinically-Informed Vision–Language Modeling Framework for Detecting Pro-Bigorexia Videos on TikTok
Minh Duc Chu1, Kshitij Pawar1, Zihao He1, Roxanna Sharifi2, Ross Sonnenblick3, Magdalayna Curry4, Laura D’Adamo3, Lindsay Young4, Stuart B Murray5, Kristina Lerman1 1USC Information Sciences Institute 2Keck School of Medicine, USC 3Department of Clinical Psychology, Drexel University 4Annenberg School for Communication and Journalism, USC 5Department of Psychiatry and Biobehavioral Sciences, UCLA 1mhchu@usc.edu
1 Introduction
Figure 1: BigTok pipeline overview. Left: Expert annotation process with dual annotation and consensus resolution. Right: Multimodal feature extraction (visual, audio, text) and classification evaluation across VLMs using zero-shot, few-shot, and finetuning approaches for primary category, subcategory, and severity prediction tasks.
Social media platforms face mounting challenges in detecting harmful content that significantly impacts user mental health Chancellor and Choudhury (2020); Gorwa et al. (2020). Contemporary platforms like TikTok algorithmically amplify content promoting unrealistic body ideals Becker (2004); Minadeo and Pope (2022), yet much harmful material exists in gray areas where legitimate discussions intersect with dangerous messaging Gillespie (2018). These narratives emerge through complex multimodal combinations—visual imagery, audio cues, and textual descriptions—that traditional text-based detection systems cannot adequately capture Kiela et al. (2020a, b). Effective automated moderation requires sophisticated multimodal approaches that can parse nuanced signals across video, audio, and text Gimeno-Gómez et al. (2024).
Pro-bigorexia content—material promoting muscle dysmorphic behaviors—exemplifies these multimodal detection challenges at their most complex. This harmful content promotes compulsive muscle-building behaviors that disproportionately affect adolescent males Pope Jr et al. (1997); Mitchison et al. (2022), yet remains severely understudied in computational research compared to traditional eating disorder detection Chancellor and Choudhury (2020). Pro-bigorexia material masquerades as legitimate fitness content through subtle combinations of muscular displays, extreme workout demonstrations, and coded supplement language Murray et al. (2017); Kamkari (2025). The visual similarity between harmful and beneficial fitness content, combined with evolving coded terminology, renders current text-based detection approaches fundamentally inadequate.
Current automated detection systems face critical limitations when confronting such content. Foundation models trained on general corpora lack domain-specific knowledge to recognize subtle clinical markers and euphemistic language patterns Murray et al. (2017). Existing systems rely heavily on keyword filtering and text-based signals, missing critical visual and behavioral cues embedded in video content Kiela et al. (2020b); Gorwa et al. (2020). The dynamic nature of social media—where creators deliberately evolve language to evade detection—further challenges automated approaches Gillespie (2018). Effective detection requires integrated analysis of visual behavioral cues, audio transcripts, and textual context—capabilities that remain underexplored in computational mental health research Gimeno-Gómez et al. (2024).
To address these detection challenges, we make the following contributions:
  • We develop the first clinically-informed taxonomy for automated pro-bigorexia detection, establishing fine-grained categories spanning body image, nutrition, supplement abuse, exercise practices, and masculinity that enable systematic computational analysis.
  • We introduce BigTok, a multimodal dataset of over 2,200 expert-annotated TikTok videos that enables robust evaluation of vision-language models on challenging harmful content detection tasks.
  • We construct BigTokDetect, a detection framework that achieves state-of-the-art (SOTA) performance using proprietary and open-source VLMs, with our best models reaching 95.1% accuracy on primary category classification and 91.0% on fine-grained subcategory detection through fine-tuning.
  • We demonstrate through comprehensive ablation studies that multimodal fusion improves detection performance by 5-10% over text-only approaches, with video features providing the most discriminative signals for identifying subtle pro-bigorexia behaviors.
Figure 1 summarizes our comprehensive approach to multimodal pro-bigorexia detection. Our methodology demonstrates that clinically-informed vision-language models can effectively identify subtle harmful content that evades traditional detection systems. Through systematic evaluation across multiple model architectures and training paradigms, we achieve substantial improvements over existing text-based methods, establishing a new paradigm for detecting nuanced harmful content on social media platforms. The BigTokDetect framework provides both the computational tools and replicable methodology needed to address the growing challenge of automated content moderation in mental health domains.
2 Related Work
Muscle Dysmorphic Disorder on Social Media
Muscle dysmorphic disorder, or "bigorexia," involves preoccupation with insufficient muscularity, driving compulsive behaviors including exercise, rigid dieting, and supplement abuse Pope Jr et al. (1997); Cooper et al. (2020). Frequent exposure to muscularity-oriented TikTok and Instagram content—including body-transformation videos, supplement promotions, and steroid use—has been linked to elevated rates of probable muscle dysmorphia Mitchison et al. (2022); Ganson et al. (2025). Hypermasculine online subcultures (e.g., the "manosphere") intensify these pressures through toxic social comparisons and body-optimization trends Kamkari (2025). Despite this prevalence, both psychological research and platform moderation remain focused on anorexia and bulimia, leaving bigorexia largely unmoderated Lookingbill et al. (2023).
Multimodal Mental Health and Body Image Content Detection
Automated detection of body image content has overwhelmingly targeted anorexia and bulimia Chancellor et al. (2017); Chancellor and Choudhury (2020), relying primarily on keyword filters and text-only classifiers despite evidence that vision-language fusion substantially improves detection of subtle harmful imagery Kiela et al. (2020b). Short-form video platforms have become dominant for youth mental health discourse Basch et al. (2022), yet eating disorder detection remains predominantly text-based Wang et al. (2017); Merhbene et al. (2024). Even dedicated TikTok corpora for eating disorders research exclude muscle dysmorphia labels Bickham et al. (2025); Donati et al. (2023), while existing qualitative analyses of pro-muscularity forums Murray et al. (2015) lack the scale needed for robust computational detection. Modern vision-language models—Flamingo Alayrac et al. (2022), InstructBLIP Li et al. (2023), GPT-4V OpenAI (2023)—enable end-to-end multimodal detection but lack clinically grounded training data for specialized tasks. Existing mental health datasets rely on crowdsourced annotations rather than clinical expertise, potentially missing subtle markers needed to distinguish legitimate fitness content from harmful pro-bigorexia messaging that uses coded language around supplements and steroids Murray et al. (2017). To our knowledge, no large-scale expert-annotated datasets exist for this increasingly prevalent content type.
3 Pro-bigorexia Content Taxonomy
We develop an empirically grounded taxonomy for pro-bigorexia content characterization, following an iterative four-step process. First, we compile a seed list of keywords from fitness, bodybuilding, exercise, and diet domains, supplemented by terms from foundational muscle dysmorphia research and popular media reports Cafri et al. (2008); Sánchez et al. (2024); Willett (2023). Then, using these keywords, we sample 500 TikTok videos over 40 hours, and open-code emergent patterns of bigorexia, such as body-check demonstrations, supplement endorsements, extreme workout narratives, diet tutorials). Next, we draft preliminary categories based on the DSM-5 clinical criteria American Psychiatric Association (2013), validated measures such as the Drive for Muscularity Scale McCreary and Sasse (2000) and the Muscle Dysmorphic Disorder Inventory Zeeck et al. (2018). Finally, we convene with clinical experts specializing in eating disorders and body image psychopathologies to refine categories, clarify specialized or coded language (especially around steroids and supplements), and disambiguate edge cases. We iterate to discover new potential categories and keywords.
Our final taxonomy for pro-bigorexia TikTok content consists of five primary categories: Relationship to Body, Relationship to Food, Relationship to Exercise, Supplements, and Masculinity. Each is further divided into clinically grounded subcategories (Table 12, Appendix A.1). For each category and subcategory, we provide clear definitions, representative keywords, practical guidelines for recognizing problematic content, and sample videos. This approach yields a consistent, multimodal annotation of overt and implicit signals of pro-bigorexia behaviors.
4 BigTok: Expert-Curated Multimodal Dataset Construction
4.1 Data Collection
We source our corpus from TikTok via its official API TikTok (2025). We query videos from January 2019 to January 2025 to capture pre- and post-pandemic trends. Guided by domain experts, we curate 40 high-precision query terms mapped to taxonomy subcategories (Table 12, Appendix A.1), retrieving up to 1,000 videos and their metadata per term. We randomly sample videos per group of keywords that belong to each primary and subcategory to 2,400 videos for annotation. Only video content and captions are exposed to annotators to protect user privacy (anonymization steps are described in Appendix C). We also compile negative (irrelevant) examples using 42 trending hashtags from TikTok’s Creative Center TikTok Creative Center , following established supervised classification practices Kiela et al. (2020b).
4.2 Annotating Bigorexia Content
Annotator Recruitment and Demographics
We recruit 16 subject matter experts (13 females and 3 males), including licensed clinical psychologists, social workers, and doctoral candidates with research and clinical experience in eating and body image disorders (Table 11, Appendix B.1 for profiles). Their combined expertise in empirical research and direct patient care ensures that our annotations are both theoretically grounded and have clinical relevance.
Annotation Instructions
We base our annotation interface on Amazon Mechanical Turk via an invitation-only pool (Appendix B.3). Annotators are instructed to select the first (mandatory) primary–subcategory pair ($t_{1},s_{1}]$) and the second (optional) primary–subcategory pair ($t_{2},s_{2}]$). Every video is also rated for harm on a discrete 5-point Likert scale with 0.5-point increments (1=not harmful and 5=very harmful). The annotation interface is shown in Appendix B.2.
Annotator Training
As a pilot study, each annotator independently labels a set of 20-30 videos to validate taxonomy coverage, identify potential edge cases, and familiarize themselves with the annotation interface. We monitor disagreements and encourage detailed note-taking. Annotators then join a group session to discuss flagged issues, collaboratively annotate selected videos while explaining their reasoning, until consensus is reached.
Batchwise Annotation
We split the dataset into 8 batches of ~300 videos. For each batch, we filter out videos that are not available (e.g., they have been deleted) or are not in English. Then the batch is assigned to two different annotators, so that every video receives two independent annotations. To assess agreement between a pair of annotators A and B, we compare all combinations of labels and determine the level of agreement:
  • Perfect agreement: both the first and second primary–subcategory pairs match exactly across annotators, i.e. $\forall i\in\{1,2\}:\;[t_{i},s_{i}]^{A}=[t_{i},s_{i}]^{B}$
  • Strong agreement: one primary–subcategory pair matches exactly across annotators, i.e. $\exists i\in\{1,2\}:\;[t_{i},s_{i}]^{A}=[t_{i},s_{i}]^{B}$
  • Weak agreement: at least one primary category matches across any labels, but the subcategories differ, i.e. $\exists i\in\{1,2\}:\;t_{i}^{A}=t_{i}^{B}\;\wedge\;s_{i}^{A}\neq s_{i}^{B}$
  • Disagreement: no common primary category across labels, i.e. $\forall i\in\{1,2\},\;t_{i}^{A}\neq t_{i}^{B}\;\wedge\;s_{i}^{A}\neq s_{i}^{B}$.
We then flag examples with weak agreement or disagreement on pro-bigorexia detection, and examples where severity scores differ by more than two points. The two annotators are invited to a structured discussion to articulate their reasoning, review cases that need consultation, address disagreements, and eventually aim to reach consensus, although consensus is not mandatory. After the discussion, each annotator independently re-annotates the flagged videos. Further details on this procedure are provided in Appendix  B.4.
Label Aggregation
While each video may receive multiple label pairs (primary-subcategory), for sake of easy evaluation, we assign a single label pair to each video. For videos with perfect agreement, we randomly choose of the label pairs. For strong agreement, we adopt the shared primary–subcategory pair; for weak agreement, we choose the mandatory (first) label over the optional (second) one; remaining disagreements are resolved by random tie-breaking. The final harm severity score is the average of the two annotators’ ratings. This simplification reflects current limitations of LLMs in reliable multi-label classification Ma et al. (2025).
We assess inter-rater reliability with Cohen’s $\kappa$  Cohen (1960) for two annotators. Initial agreement is moderate (strong $\kappa=0.43$$0.69$; weak $\kappa=0.59$$0.83$; Table 6). After structured consensus discussions and re-annotation, reliability increases substantially (strong $\kappa=0.58$$0.81$; weak $\kappa=0.78$$0.94$), ensuring that our iterative process yields high-quality, consistent labels. All Cohen’s $\kappa$ values for each batch are shown in Table 6 (Appendix B.5)
4.3 Data Statistics
Primary Category Count
Relationship to Body 449
Relationship to Exercise 450
Relationship to Food 317
Supplement Abuse 220
Relationship to Masculinity 131
Irrelevant 639
Subcategory Count
Muscularity Self-objectification 278
Leanness Self-objectification 72
Muscle Dissatisfaction 78
Rigid Food Rules 136
Unsafe Food 79
Cheat Meals 74
Excessive Exercise 109
Predebting Exercise 22
Maladaptive Coping 89
Exercise-Induced Functional Impairment 34
Toxic Motivation 107
Anabolic Steroids 68
Legal APEDs 50
Hormone Therapy 82
Relationship to Masculinity 131
Other 103
Irrelevant 639
Table 1: Distribution of categories and subcategories BigTok. We include the primary categories Irrelevant and Relationship to Masculinity, which can also be considered as subcategories.
Our final BigTok dataset comprises 2,210 annotated TikTok videos. The median video duration is 20.6 seconds, and 59.8% of videos are under 30 seconds. Table 1 shows the label distribution. At the primary category level, excluding Irrelevant, Relationship to Exercise (450) is the most dominant class. At the subcategory level, Muscularity Self-objectification predominates (278), reflecting the popularity of the genre of videos in which people display their muscles. The Other subcategory covers videos within a primary category that do not match predefined subcategories (e.g., humor or lifestyle clips); these are excluded from subcategory-level benchmarking. Some examples from BigTok are shown in Tables 13, 14, and 15 Appendix A.1. Additional annotation statistics are provided in Appendix B.5.
5 Classifying Multimodal Pro-Bigorexia Content via VLMs
To develop an automated detection system for pro-bigorexia content, we apply SOTA VLMs to this challenging multimodal classification task. We leverage both commercial and open-source VLMs through zero-shot prompting, few-shot learning, and fine-tuning approaches to achieve high performance in identifying subtle pro-bigorexia behaviors across video, audio, and text modalities.
5.1 Task Definition
We define three evaluation tasks for TikTok video classification. Task 1: Primary Category Classification involves predicting one of the five primary categories of pro-bigorexia content (Body, Food, Supplements, Exercise, Masculinity), or Irrelevant. Task 2: Subcategory Classification involves predicting the specific subcategory within the selected primary category. Task 3: Severity Estimation involves predicting a continuous severity score on a 1-5 scale, where 1 indicates no harm and 5 indicates extreme harm.
For each task, we split the data into training and test sets at a 3:1 ratio using stratified sampling. The test set is strictly balanced across all classes, while the training set is adjusted by downsampling the majority categories to mitigate class imbalance. The sampling procedures and statistics after sampling are provided in Appendix C.
5.2 Models
Given the multimodal nature of our classification task, involving text from captions, audio transcripts, images, and video content, we require SOTA models designed for multimodal data, particularly video understanding. We evaluate three commercial API-based VLMs: GPT-4.1 OpenAI (2025) and Claude-Sonnet-4 Anthropic PBC (2025) (both process images), while Gemini-2.5-Flash Comanici et al. (2025) natively accepts video input. We also evaluate two open-source models: Qwen2.5-VL Wang et al. (2024), and InternVL3 Zhu et al. (2025), which uses Qwen2.5 pre-trained base models and Variable Visual Position, which achieves strong performance on video benchmarks.
Model sizes and versions are listed in Table 10, Appendix D.1. Due to computational constraints, we focus on small and medium-sized models. For open-source models, we set temperature = 0.1 and top_p = 0.9. For API models, we retain the default hyperparameters temperature = 1.0 and top_p = 1.0. Implementation details and hardware specifications are provided in Appendix D.1. Details about the train and test datasets are in Section C, Appendix.
5.3 Input Features
To capture the full range of pro-bigorexia signals, we extract four complementary input features: Visual: For models that natively accept video (e.g., Gemini-2.5-Flash, Qwen2.5VL, InternVL3), we provide raw video. For frame-based models (e.g., Claude-Sonnet-4, GPT-4.1, open-source VLMs), we sample four equally spaced frames per video due to API costs and token-length constraints; ablation studies (Section 5.7) show additional frames yield minimal gains. Audio: We apply Google’s YAMNet Hershey et al. (2017) to filter non-speech sounds (music, ambient noise), then transcribe "Speech" segments using OpenAI Whisper Radford et al. (2022). On-Screen Text: We use Gemini-2.5-Flash to extract overlaid text—captions and annotations—since many TikTok creators rely on on-screen text for key messages. Caption: We include the original TikTok description (user caption and hashtags) from the video metadata.
Model Training Primary Category Subcategory
Acc. Pm Rm F1m Acc. Pm Rm F1m
GPT-4.1 Zero-shot 0.796 0.808 0.792 0.792 0.652 0.639 0.503 0.532
Few-shot 0.813 0.820 0.813 0.813 0.680 0.679 0.643 0.639
Claude-Sonnet-4 Zero-shot 0.829 0.832 0.829 0.827 0.670 0.679 0.636 0.640
Few-shot 0.819 0.829 0.818 0.818 0.665 0.674 0.627 0.632
Gemini-2.5-Flash Zero-shot 0.776 0.782 0.776 0.775 0.666 0.665 0.616 0.621
Few-shot 0.776 0.782 0.776 0.775 0.663 0.645 0.613 0.621
Qwen2.5-VL-7B Zero-shot 0.539 0.744 0.539 0.526 0.383 0.554 0.327 0.306
Few-shot 0.689 0.733 0.689 0.689 0.236 0.149 0.142 0.140
Finetuning 0.784 0.810 0.784 0.776 0.684 0.687 0.669 0.667
Qwen2.5-VL-32B Zero-shot 0.733 0.788 0.733 0.742 0.556 0.637 0.495 0.513
Few-shot 0.804 0.812 0.804 0.807 0.654 0.662 0.594 0.607
Finetuning 0.776 0.800 0.776 0.775 0.658 0.662 0.647 0.645
InternVL3-8B Zero-shot 0.614 0.734 0.614 0.635 0.517 0.582 0.464 0.476
Few-shot 0.690 0.717 0.690 0.692 0.519 0.560 0.436 0.443
Finetuning 0.765 0.789 0.765 0.758 0.690 0.686 0.679 0.675
InternVL3-38B Zero-shot 0.784 0.797 0.784 0.785 0.649 0.665 0.603 0.608
Few-shot 0.806 0.817 0.806 0.807 0.673 0.666 0.630 0.630
Finetuning 0.767 0.795 0.767 0.758 0.636 0.540 0.535 0.528
Table 2: Classification Macro Metrics for Primary Category and Subcategory by Model and Training Method. Highest values in each column are in green, second-highest in cyan. Pm, Rm, and F1m refer to macro precision, macro recall, and macro F1 scores.
5.3.1 Training and Inference Paradigms
Zero-Shot Prompting
We create prompts for zero-shot classification by incorporating taxonomy definitions of the primary and subcategories and instructing the model to select the appropriate category as the label. The full zeroshot prompt templates for Tasks 1 and 2 are shown in Figure 6 and 7, Appendix D.2. We evaluate the models on balanced test sets of 588 examples for primary classification (Task 1) and 466 examples for subcategory classification (Task 2), with each class evenly represented (see Appendix C)
Few-Shot Prompting
For few-shot experiments, we sample a fixed set of in-context examples from the training data: two videos per primary category (12 examples total) for Task 1 and one video per subcategory (16 examples) for Task 2 (prompt details are in Figures 8 and 9, Appendix). Each example is presented with its full multimodal features: video frames, audio transcript, on-screen text, caption, and the corresponding label. We evaluate the VLMs on a balanced test set of 462 examples (Appendix C).
Finetuning
Due to cost constraints, we limit finetuning to open-source VLMs. We instruction-tune various model variants (up to 72B) on our annotated datasets. The batch size is 8, and the learning rate is 5e-5. The inference hyperparameters are similar to zero-shot and few-shot prompting (temperature = 0.1 and top_p = 0.9).
5.4 Pro-Bigorexia Classification Results
Primary Category Detection
Commercial API-based models continue to lead in broad-category detection: Claude-Sonnet-4 (zero-shot) attains the highest accuracy (0.829), macro precision (0.832), macro recall (0.829), and macro F1 (0.827), with its few-shot variant ranking a close second. Although fine-tuned open-source models (e.g., Qwen2.5-VL-32B and InternVL3 variants) gain 10–20 points over their zero-shot baselines, none yet surpass the commercial zero-shot benchmark. These results confirm that large-scale pretraining still confers an advantage for general category classification, even as domain-specific tuning significantly narrows the gap.
Subcategory Detection
In contrast, fine-tuning proves decisive for fine-grained subcategory tasks. InternVL3-8B (finetuned) achieves the top subcategory accuracy (0.690), recall (0.679), and F1 (0.675), while Qwen2.5-VL-7B (finetuned) leads in precision (0.687) and ranks second in other metrics. Commercial zero- and few-shot methods plateau around 0.670 accuracy and underperform relative to these tuned open-source variants. This demonstrates that parameter adaptation is essential to capture the nuanced distinctions of harmful content subtypes.
Prompting vs. Parameter-Tuning
Few-shot prompting yields modest boosts in primary-category performance—for instance, GPT-4.1 improves from 0.796 to 0.813 accuracy—highlighting the continued value of emergent prompting abilities. However, in subcategory detection, few-shot gains are minimal, and zero-shot performance remains low. By contrast, parameter-efficient fine-tuning delivers substantial improvements across both tasks, underscoring that direct model adaptation is the most effective strategy for specialized harmful content detection, particularly for granular subcategory classification.
Model Scale and Architecture
Larger Qwen2.5-VL-32B boosts zero/few-shot primary category performance, but mid-scale InternVL3-8B (built on the Qwen backbone) achieves the strongest fine-tuning gains—surpassing its 38B variant on subcategory tasks. This underscores that, beyond raw parameter count, architecture critically shapes fine-grained detection performance.
5.5 Severity Score Prediction
Model Severity
MAE $\downarrow$ $\rho$ $\uparrow$
Claude-Sonnet-4 0.679 0.675
GPT-4.1 0.690 0.690
Gemini-2.5-Flash 0.690 0.690
InternVL3-8B 0.805 0.474
InternVL3-38B 0.701 0.607
Qwen2.5-VL-7B 0.794 0.484
Qwen2.5-VL-32B 0.688 0.601
Qwen2.5-VL-72B 0.653 0.638
Table 3: Zero-shot Severity Prediction Metrics by Model (highest in green, 2nd highest in blue).
To assess the ability of VLMs to predict harm severity, we conducted zero-shot prompting experiments across multiple models. Results reveal that models typically misestimate severity by approximately half a scale point, indicating reasonable but imperfect alignment with expert clinical judgment. Notably, severity prediction proves more challenging than categorical classification, with all models showing only moderate correlation with ground truth ratings. While these findings demonstrate baseline competence in severity estimation, substantial improvement is needed for reliable moderation deployment.
5.6 Error Analysis
Cross-model error analysis exposes fundamental challenges in characterizing pro-bigorexia content. For instance, models consistently misclassify exercise-related content as “Irrelevant” (6.1–14.3%) or confuse “Relationship to Body” with “Relationship to Exercise” (10.2–17.3%), mirroring annotator disagreements (Figures 3 and 4, Appendix B.5). Videos demonstrating workout sessions simultaneously touch on exercise, body display, dieting, lifestyle, and motivations that single-label classification cannot capture.
Classification consistently confuses “Hormone Therapy” and “Anabolic Steroids” (11.8–17.6% bidirectional misclassification), a clinically critical distinction. Models also frequently misclassify “Supplement Abuse” as “Relationship to Body” (8.2–22.4%), suggesting difficulty in recognizing harmful supplement messaging. This likely stems from creators using coded language (e.g., “tren”, “stack”), requiring domain expertise that general-purpose models lack. Additionally, models show systematic bias toward predicting “Muscularity Self-objectification” in subtype classification, with excessive exercise and toxic motivation content frequently misclassified into this category (14.7–47.1%). These systematic errors highlight a fundamental challenge: effective pro-bigorexia detection demands not just multimodal capabilities, but clinical and social knowledge to navigate the blurred boundaries between fitness content and promoting unhealthy behaviors.
5.7 Ablation Study
5.7.1 Input Features: Text vs Video
Model Modality Pm Rm F1m
Gemini-2.5-Flash Audio 0.700 0.423 0.424
Caption 0.732 0.707 0.709
OCR 0.717 0.661 0.664
Text 0.733 0.719 0.719
Video 0.786 0.765 0.766
All 0.782 0.776 0.775
GPT-4.1 Audio 0.752 0.444 0.453
Caption 0.728 0.672 0.679
OCR 0.741 0.663 0.672
Text 0.749 0.675 0.681
Video 0.755 0.697 0.701
All 0.808 0.792 0.792
InternVL3-38B Audio 0.632 0.405 0.416
Caption 0.669 0.590 0.587
OCR 0.649 0.614 0.618
Text 0.691 0.630 0.635
Video 0.697 0.655 0.660
All 0.797 0.784 0.785
Table 4: Ablation results for task 1 using a single modality input features (text or video). Audio modality refers to the transcript of the video; Caption refers to the video description; OCR is text within images, and Text modality refers to Caption + OCR. Highest values in each column are in green, second-best in cyan. Pm, Rm, and F1m refer to macro precision, macro recall, and macro F1 scores.
Table 4 reports results of ablation study to assess the contribution of each modality to classification performance on Task 1 with zero-shot prompting. We evaluate GPT-4.1 and InternVL3-38B (image-based) and Gemini-2.5-Flash (video-native). Multimodal fusion yields the best results, with GPT-4.1 and InternVL3-38B achieving F1 scores of 0.792 and 0.785, respectively. Video alone provides the strongest unimodal performance (F1 = 0.660–0.701), highlighting the discriminative power of visual behavioral cues. In contrast, audio consistently underperforms (F1 << 0.453), likely due to background music and variable audio quality. Gemini-2.5-Flash shows narrower modality gaps, suggesting native video processing reduces the need for explicit fusion. Overall, results underscore the value of combining visual and textual inputs for robust pro-bigorexia detection.
5.7.2 Number of Frames per Video
Model Frames # Pm Rm F1m
GPT-4.1 4 0.832 0.829 0.827
16 0.817 0.793 0.792
32 0.739 0.657 0.685
InternVL3-38B 4 0.7967 0.7840 0.7849
16 0.7988 0.7891 0.7896
32 0.7922 0.7840 0.7844
Table 5: Frame-count ablation for GPT-4.1 and InternVL3-38B. Pm, Rm, and F1m are macro precision, macro recall, and macro F1.
[…]
6 Conclusion
We present BigTok and BigTokDetect, a clinically-informed dataset and vision–language framework for identifying pro-bigorexia content on TikTok. By curating over 2,200 expert-annotated videos across five primary categories and fine-grained subcategories, we enable systematic study of muscle dysmorphic signals that evade traditional text-based moderation. Our fine-tuned VLMs achieve 82.9% accuracy on primary category classification and 69.0% on subcategory detection, and ablation studies confirm that multimodal fusion of visual, audio, and textual inputs yields a 5–10% boost over unimodal baselines.
Despite these advances, challenges remain in scaling to continually evolving platforms and detecting emerging coded language. Future work expands BigTok to include cross-platform content (e.g., Instagram, YouTube Shorts), incorporates temporal dynamics of video sequences, and integrates user-level metadata for context-aware moderation. We also collaborate with clinical practitioners to validate model outputs in real-world settings and explore fairness considerations across demographic groups. By releasing our dataset, code, and model checkpoints, we aim to catalyze further research on robust, ethically grounded multimodal moderation in specialized mental health domains.
Limitations
Dataset and Sampling Constraints
Our dataset is limited to English-only TikTok content from 2019-2025, potentially missing cultural variations and platform-specific differences. We focused on English to ensure annotation quality and TikTok as the dominant youth platform. Keyword-based sampling may overlook subtle or emerging pro-bigorexia content that avoids our taxonomy terms, though we exhaustively developed keywords through previous literature review and expert consultation to maximize coverage.
Annotation and Taxonomy Limitations
Forced reduction from multi-faceted expert annotations to single labels loses nuanced information. We simplified to single labels due to current LLM limitations in reliable multi-label classification. Our 16 annotators are predominantly female (13/16), potentially introducing gender perspective bias in male-centric bigorexia evaluation; however, this field remains understudied, and we could not find sufficient male experts specializing in bigorexia. Our taxonomy represents a living document that may miss emerging patterns, though we strived for exhaustiveness through iterative expert refinement.
Model Selection Limitations
While we selected current leading VLMs across both commercial and open-source categories, the rapidly evolving landscape means we may be missing other capable models. We prioritized models with proven multimodal video understanding capabilities, but acknowledge that newer or specialized architectures might offer different performance characteristics.
Experimental Reproducibility
Due to computational cost constraints, we conduct single runs for each experimental configuration without multiple trials or statistical significance testing. While our results establish baseline performance across models and tasks, future work should include multiple experimental runs with statistical analysis to provide more robust performance comparisons and confidence intervals for the reported metrics.
Ethics Statement
This research addresses harmful content classification, which raises important ethical considerations. To protect annotators’ well-being when reviewing potentially disturbing pro-bigorexia content, we exclusively recruited clinical experts—licensed psychologists, psychiatrists, and doctoral candidates specializing in eating disorders—who possess the professional training necessary to safely engage with this sensitive material. Their clinical expertise provides essential psychological safeguards while ensuring high-quality annotations. While our work aims to support mental health research, we acknowledge that automated detection systems risk censorship or discrimination against legitimate fitness communities. We took precautions, including content anonymization and clinical expert consultation, focusing on detection research rather than deployment recommendations. Our work is intended for research purposes and should not be used for clinical diagnosis of bigorexia without proper clinical oversight. All procedures were conducted in accordance with institutional review board guidelines.
We acknowledge the use of AI language models to assist with writing and editing portions of this manuscript. All AI-generated content was reviewed, verified, and edited by the authors. The research design, data collection, analysis, and primary contributions remain entirely the work of the human authors.
Appendix
Appendix A Taxonomy
A.1 Categories Definition
A detailed definition of the categories in our taxonomy is provided in Table 12. These categories are carefully curated and iteratively refined in collaboration with clinical psychologists and domain researchers to ensure both validity and practicality.
Appendix B Annotation
B.1 Annotator Profiles
We enlist 16 subject-matter experts, ranging from clinical psychologists and psychiatrists to computational social scientists, for video annotation. Our annotators are research collaborators and co-authors who volunteered their clinical expertise with full knowledge of the project’s scope and scientific objectives, ensuring informed participation in this sensitive content annotation task. Table 11 details each annotator’s ID, batch assignment, area of expertise, and gender.
B.2 Annotation Instructions
Annotators received comprehensive guidelines emphasizing safety protocols and decision-making frameworks for handling potentially disturbing content. Key instructions included: (1) Content Safety: annotators were advised to work at their own pace, take regular breaks, and prioritize mental well-being when reviewing harmful content; (2) Decision Framework: when encountering ambiguous cases, annotators were instructed to select “Unsure, need consultation” and provide explanatory notes, with textual information (captions, on-screen text) prioritized over conflicting visual content; (3) Consistency Maintenance: annotators were asked to reflect on their annotation patterns periodically to avoid drift and maintain consistency across the 300-video batches; (4) Expert Support: availability of research coordinators via dedicated communication channels for immediate consultation on challenging cases.
Annotators assess harm severity based on message intensity, including explicit display of muscular physiques, toxic motivational pressure toward unattainable muscularity, rigid and unrealistic dietary demands, endorsement of high-risk behaviors such as steroid use or dangerous workout practices, and the degree of pathological behavioral patterns exhibited in the content.
These protocols ensured both annotator safety and annotation quality while acknowledging the inherently difficult nature of distinguishing harmful pro-bigorexia content from benign fitness material.
B.3 Annotation Platform
The HTML interface (Figure 2) presents each TikTok video with its caption and hashtags (if any). The form beneath the video prompts the annotator to mark if the video is: (1) Showing One’s Body (yes/no); includes (2) Commercial Sponsorship (yes/no); (3) 2 boxes for Type of harm (1 mandatory, 1 optional), consisting of primary categories and subcategories, plus “Relevant but Not Listed,” “Irrelevant,” and “Unsure—Needs Consultation” options; and (4) Severity of Harm on a discrete scale ranging from 1 (Not Harmful) to 5 (Extremely Harmful). A text box allows annotators to write additional comments and edge-case observations.
Figure 2: Annotation Interface on Amazon Mechanical Turk platform.
B.4 Annotation Processing
After Round 2, we apply a five-step pipeline to prepare the data for consensus review and final curation:
  1. 1.
    Annotation Completeness: verify that each video has exactly two independent annotations.
  2. 2.
    Data Filtering: remove videos that are unavailable (deleted) or non-English, as indicated in annotators’ notes.
  3. 3.
    Reliability Assessment: compute Cohen’s $\kappa$ for both strong (exact primary–secondary match) and weak (primary-only match) agreement to quantify inter-rater reliability.
  4. 4.
    Category Refinement Candidates: collect all videos flagged as “Unsure—Needs Consultation” or “Relevant but Not Listed.” For these, we consult with the annotators to either reassign them to existing categories, extend the taxonomy to cover frequently emerging content types, or mark them as irrelevant.
  5. 5.
    Consensus Review and Re-annotation: filter out all items by agreement level (strong, weak, disagreement), sample a subset of weak and disagreement cases for each annotator pair, hold a batch-specific meeting to discuss and jointly re-annotate selected cases to reach consensus, then return the remaining weak and disagreement items to annotators for independent re-annotation.
This pipeline ensures data consistency, highlights contentious cases for Round 3 consensus sessions, and informs any necessary taxonomy updates.
B.5 Annotation Statistics
Inter-annotator agreement substantially improves from Round 2 to Final, with Cohen’s $\kappa$ values increasing from moderate agreement (0.43-0.69 strong $\kappa$, 0.59-0.83 weak $\kappa$) to good-to-excellent agreement (0.58-0.81 strong $\kappa$, 0.78-0.94 weak $\kappa$). The consistently high weak $\kappa$ values (>0.9 in most batches) in the Final round indicate that annotators achieved near-perfect agreement on the broader annotation categories.
Batch Round 2 Final
Strong $k$ Weak $k$ Strong $k$ Weak $k$
1 0.551 0.613 0.616 0.775
2 0.570 0.588 0.762 0.841
3 0.510 0.612 0.680 0.815
4 0.551 0.785 0.811 0.937
5 0.563 0.685 0.787 0.924
6 0.555 0.723 0.742 0.938
7 0.430 0.739 0.577 0.931
8 0.685 0.825 0.769 0.938
Table 6: Cohen’s kappa coefficients for the inter-agreement rate among annotators in Rounds 2 and 3
Primary Category Count
Relationship to Body 1442
Relationship to Exercise 1233
Relationship to Food 824
Supplement Abuse 558
Relevant but Not Harmful 411
Relationship to Masculinity 406
Irrelevant 401
(a) Primary Categories
Subcategory Count
Muscularity Self-objectification 905
Leanness Self-objectification 238
Muscle Dissatisfaction 240
Rigid Food Rules 375
Cheat Meals 193
Excessive Exercise 379
Maladaptive Coping 213
Toxic Motivation 308
Anabolic Steroids 195
Other (please specify in Note) 217
(b) Subcategories (reordered)
Table 7: Distribution of Annotated Primary Categories and Subcategories
Appendix C Data
We collected videos through the official TikTok Research API after submitting a research proposal that was reviewed and approved by TikTok. This API accesses publicly posted content where users have consented to public visibility through TikTok’s terms of service. We implemented additional privacy protections by anonymizing user identifiers and exposing only video content and captions to annotators.
Our dataset comprises TikTok videos annotated for hierarchical classification across two levels of granularity. The primary category classification categorizes content into six broad categories related to body image and health behaviors, while the subcategory classification provides fine-grained classification into 16 specific subcategories. Severity score scale define the intensity of harm emotion. The severity scores are most heavily concentrated in the 1.0 to 1.5 range (Figure 5), indicating that lower severity levels are the most common. Frequency then gradually decreases as scores increase, with relatively few cases above 4.0. This suggests that mild severity is predominant in the dataset. We remove corrupted files (e.g., missing metadata) and non-English videos during preprocessing. The dataset is split into train/test sets, maintaining approximately a 3:1 ratio with stratified sampling based on task labels and downsampling of dominant categories to improve balance.
As shown in Table 8, the primary category task contains 1,966 training and 588 test samples. Training data exhibits natural class imbalance reflecting real-world distribution, ranging from 489 samples (“Relationship to Body”) to 66 samples (“Relationship to Masculinity”), while the test set maintains balanced representation (98 samples per category) for fair evaluation. The subcategory task operates on a filtered subset of 1,472 training and 462 test samples (Table 9). Training samples range from 200 (“Irrelevant”) to 28 (“Predebting Exercise”), while test data maintains balanced distribution for major categories (34 samples each) with reduced representation for rare categories (4-10 samples).
Zero-shot and few-shot approaches are evaluated solely on test sets, while finetuning models are trained on the respective task’s training set and evaluated on the corresponding test set.
Primary Category Train Test
Relationship to Body 489 98
Irrelevant 450 98
Relationship to Exercise 450 98
Relationship to Food 310 98
Supplement Abuse 201 98
Relationship to Masculinity 66 98
Total 1,966 588
Table 8: Distribution of samples across primary categories (Task 1).
Subcategory Train Test
Muscularity Self-objectification 170 34
Leanness Self-objectification 59 34
Muscle Dissatisfaction 72 34
Rigid Food Rules 144 34
Unsafe Food 59 34
Cheat Meals 60 34
Excessive Exercise 129 34
Predebting Exercise 28 4
Maladaptive Coping 82 34
Exercise-Induced Functional Impairment 43 6
Toxic Motivation 102 34
Anabolic Steroids 61 34
Legal APEDs 56 10
Hormone Therapy 77 34
Relationship to Masculinity 130 34
Irrelevant 200 34
Total 1,472 462
Table 9: Distribution of samples across subcategories (Task 2).
Figure 3: Inter-annotator agreement matrix for primary categories. Each cell shows the number of videos where Annotator A assigned the row category and Annotator B assigned the column category. Diagonal elements represent perfect agreement, while off-diagonal elements indicate disagreements between annotators.
Figure 4: Inter-annotator agreement matrix for subcategories. Each cell shows the number of videos where Annotator A assigned the row subcategory and Annotator B assigned the column subcategory. The matrix reveals patterns of confusion between semantically related subcategories and overall annotation consistency.
Figure 5: Distribution of Severity Scores. The histogram shows most values concentrated between 1.0 and 2.0, with fewer cases at higher severity levels.
Appendix D Modeling
D.1 Experiment Setup
We implement the VLMs using Transformers-based (for InternVL3) and vLLM Kwon et al. (2023) (for Qwen-2.5VL) implementations through the LlamaFactory framework Zheng et al. (2024), which provides efficient inference and serving capabilities for VLMs. We used pre-trained VLMs according to their respective licenses and terms of use: commercial API-based models (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Flash) under their standard API terms, and open-source models (InternVL3, Qwen2.5-VL) under their permissive licenses for research use.
Due to hardware constraints, we exclude InternVL-38B and InternVL-78VL from certain experiments. Our current infrastructure is incompatible with the vLLM version required for multi-node deployment of these larger models. Since these models exceed the memory capacity available on single nodes in our cluster, we cannot accommodate their substantial memory requirements within our computational environment.
We consider conducting experiments with LLaVA-NeXT-Video Liu et al. (2024), which achieved SOTA open-source performance on Video-MME Fu et al. (2025). However, we exclude LLaVA-NeXT-Video from our study since its context window limitation (4,096 tokens) cannot accommodate our prompting setup (text + video/frames), which often exceeds that limit.
All zero-shot experiments on open-source VLMs are conducted on 8 $\times$ NVIDIA H100 GPUs. For few-shot and finetuning experiments, we use 32 $\times$ NVIDIA A100 GPUs for Qwen2.5-VL-7B, Qwen2.5-VL-32B, InternVL3-8B, and InternVL3-38B, and 64 $\times$ NVIDIA A100 GPUs for Qwen2.5-VL-72B.
We focus on evaluating two tasks: primary category and subcategory classification. For each task, we split the data into training and test sets at a 3:1 ratio using stratified sampling. The test set is strictly balanced across all classes, while the training set is adjusted by downsampling the majority categories to mitigate class imbalance. Detailed split statistics and sampling procedures are provided in Appendix C.
Model Version Size (B) Type
GPT-4o 2024-08-06 API
Claude Sonnet 4 2024-02-24 API
Gemini 2.5 Pro 2025-06-17 API
InternVL3 2024-10-04 8, 38 Open
Qwen2.5-VL 2025-06-05 7, 32, 72 Open
Table 10: Model version dates, parameter counts, and types (API vs. Open-source) for video-based VLMs.
D.2 Prompt Templates
We use a JSON-based prompt structure following the standard role-content format for chat-based LLMs called ‘sharegpt’. Each prompt consists of:
  • A sequence of images from the video (sampled frames)
  • A user message containing the video caption, embedded text and audio transcription
  • A single assistant response with the predicted label
An example format of prompt used for zero-shot inference/supervised finetuning and few-shot inference is shown in Figures 6 and 7 and Figure 8 and 9:
VLM Zero-Shot/SFT Prompt - Task 1: Primary Category Classification
<video>You are an expert classifier of harmful pro-muscle-dysmorphia content
on TikTok.
Taxonomy:
Type 1: Relationship to Body:
  • Subtype 1: Muscularity Self‑Objectification:Idealized imagery emphasizing
  muscular aesthetics.
  • Subtype 2: Leanness Self‑Objectification: Self‑evaluation against a lean ideal.
  • Subtype 3: Muscle Dissatisfaction: Perceived insufficient muscularity.
Type 2: Relationship to Food:
  • Subtype 1: Rigid Food Rules: Obsessive tracking, bulking, or cutting.
  • Subtype 2: Unsafe Food: Promoting unconventional muscle-enhancing foods.
  • Subtype 3: Cheat Meals: Binge-compensation cycles.
Type 3: Supplement Abuse:
  • Subtype 1: Anabolic Steroids: Downplaying risks of steroid use.
  • Subtype 2: Legal APEDs: Excessive legal supplement usage.
  • Subtype 3: Hormone Therapy: Promoting TRT as performance enhancer.
Type 4: Relationship to Exercise:
  • Subtype 1: Excessive Exercise despite harm.
  • Subtype 2: Predebting: Exercising to justify eating.
  • Subtype 3: Maladaptive Coping: Sole coping method.
  • Subtype 4: Functional Impairment.
  • Subtype 5: Toxic Motivation.
Type 5: Relationship to Masculinity:
  • Subtype 1: Linking muscles to male identity and worth.
Type 6: Irrelevant:
  • Subtype 1: Content unrelated to the above.
Carefully consider all sources of information about this video:
VIDEO_ID: <Anonymized Video ID>
Caption: <TikTok Video Caption>
Audio transcript: <Audio Transcriptions using Whisper>
Embedded text: <OCR text from video>
###TASK###
Classify the video into one of the following types:
Relationship to Body, Relationship to Food, Supplement Abuse,
Relationship to Exercise, Relationship to Masculinity, or
Irrelevant.
Only output the type label, no explanations or subtypes.
Figure 6: Prompt used for zero-shot inference and supervised finetuning for Task 1.
VLM Zero-Shot/SFT Prompt - Task 2: Subcategory Classification
<video>You are an expert classifier of harmful pro-muscle-dysmorphia content
on TikTok.
Taxonomy:
Type 1: Relationship to Body:
  • Subtype 1: Muscularity Self‑Objectification:Idealized imagery emphasizing
  muscular aesthetics.
  • Subtype 2: Leanness Self‑Objectification: Self‑evaluation against a lean ideal.
  • Subtype 3: Muscle Dissatisfaction: Perceived insufficient muscularity.
Type 2: Relationship to Food:
  • Subtype 1: Rigid Food Rules: Obsessive tracking, bulking, or cutting.
  • Subtype 2: Unsafe Food: Promoting unconventional muscle-enhancing foods.
  • Subtype 3: Cheat Meals: Binge-compensation cycles.
Type 3: Supplement Abuse:
  • Subtype 1: Anabolic Steroids: Downplaying risks of steroid use.
  • Subtype 2: Legal APEDs: Excessive legal supplement usage.
  • Subtype 3: Hormone Therapy: Promoting TRT as performance enhancer.
Type 4: Relationship to Exercise:
  • Subtype 1: Excessive Exercise despite harm.
  • Subtype 2: Predebting: Exercising to justify eating.
  • Subtype 3: Maladaptive Coping: Sole coping method.
  • Subtype 4: Functional Impairment.
  • Subtype 5: Toxic Motivation.
Type 5: Relationship to Masculinity:
  • Subtype 1: Linking muscles to male identity and worth.
Type 6: Irrelevant:
  • Subtype 1: Content unrelated to the above.
Carefully consider all sources of information about this video:
VIDEO_ID: <Anonymized Video ID>
Caption: <TikTok Video Caption>
Audio transcript: <Audio Transcriptions using Whisper>
Embedded text: <OCR text from video>
###TASK###
Classify the video into one of the following subtypes: Muscularity Self‑Objectification,
Leanness Self‑Objectification,
Muscle Dissatisfaction, Rigid Food Rules,
Unsafe Food, Cheat Meals,
Anabolic Steroids, Legal APEDs,
Hormone Therapy, Excessive Exercise,
Predebting, Maladaptive Coping,
Exercise‑Induced Functional Impairment,
Toxic Motivation,
Relationship to Masculinity, or Irrelevant.
Only output the specific subtype label, no explanations or other text.
If none apply, output Irrelevant.
Use the exact subtype names, not the type names.
Figure 7: Prompt used for zero-shot inference and supervised finetuning for Task 2.
VLM Few-Shot Prompt. Task 1: Primary Category Classification
You are an expert classifier of harmful pro-muscle-dysmorphia content on TikTok.
Taxonomy:
Type 1: Relationship to Body:
  • Subtype 1: Muscularity Self‑Objectification: Idealized imagery emphasizing muscular
  aesthetics as the primary source of value.
  • Subtype 2: Leanness Self‑Objectification: Self‑evaluation against a lean,
  low‑fat, highly‑toned ideal.
  • Subtype 3: Muscle Dissatisfaction: Expressing perceived insufficient muscularity despite
  having a muscular physique.
Type 2: Relationship to Food:
  • Subtype 1: Rigid Food Rules: Obsessive macro/micronutrient tracking
  and restrictive dietary practices to rapidly gain muscle mass (bulking) or lose fat (cutting).
  • Subtype 2: Unsafe Food: Promotion of unconventional foods believed
  to enhance muscle growth.
  • Subtype 3: Cheat Meals: Large "reward" meals after restrictive dieting
  that reinforce binge–compensation cycles.
Type 3: Supplement Abuse:
  • Subtype 1: Anabolic Steroids: Normalization or
  endorsement of anabolic‑androgenic steroid use with downplayed risks.
  • Subtype 2: Legal APEDs: Overuse of legal supplements (e.g., creatine,
  protein, pre-workout) beyond recommended doses.
  • Subtype 3: Hormone Therapy: Downplaying risks and spreading misinformation
  about testosterone replacement therapy, promoted as a performance enhancer or "anti-aging" treatment
  without proper medical diagnosis.
Type 4: Relationship to Exercise:
  • Subtype 1: Excessive Exercise: Extreme exercise routines exceeding healthy limits despite
  injury or life interference.
  • Subtype 2: Predebting: Treating exercise as punishment or permission to eat.
  • Subtype 3: Maladaptive Coping: Using exercise as the sole coping mechanism to avoid
  emotional distress.
  • Subtype 4: Exercise‑Induced Functional Impairment: Prioritizing exercise
  over essential duties, harming daily functioning.
  • Subtype 5: Toxic Motivation: Demeaning communication that pressures
  unrealistic fitness standards via shaming or slurs.
Type 5: Relationship to Masculinity:
  • Subtype 1: Linking muscle‑building and exercise performance to male identity,
  sexuality, and self‑worth.
Type 6: Irrelevant:
  • Subtype 1: Content without muscle‑obsession, restrictive
  diets, supplement/AAS promotion, or extreme exercise. Includes dance trends,
  memes, travel/cooking vlogs, general wellness (e.g., yoga), and pure entertainment unrelated
  to body‑image narratives.
Here are some examples from the training data:
Example 1:
<video>
Caption: <Video Caption>
Audio transcript: <Audio Transcription from Whisper>
Embedded text: <OCR Text>
Classification: <Type>
Example 12:
<video>
Caption: <Video Caption>
Audio transcript: <Audio Transcription from Whisper>
Embedded text: <OCR Text>
Classification: <Type>
<video>
VIDEO_ID:
Caption:
Audio transcript:
Embedded text:
###TASK###
Classify the video into one of the following types: Relationship to Body, Relationship to Food,
Supplement Abuse, Relationship to Exercise, Relationship to Masculinity, or Irrelevant.
Only output the type label, no explanations or other text. If none apply, output Irrelevant.
Don’t use a subtype, such as "Muscularity Self‑Objectification" or "Unsafe Food".
Valid outputs: Relationship to Body, Relationship to Food, Supplement Abuse, Relationship to Exercise,
Relationship to Masculinity, Irrelevant
Figure 8: Prompt used for few-shot inference for Task 1.
VLM Few-Shot Prompt. Task 2: Subcategory Classification
You are an expert classifier of harmful pro-muscle-dysmorphia content on TikTok.
Taxonomy:
Type 1: Relationship to Body:
  • Subtype 1: Muscularity Self‑Objectification: Idealized imagery emphasizing muscular
  aesthetics as the primary source of value.
  • Subtype 2: Leanness Self‑Objectification: Self‑evaluation against a lean,
  low‑fat, highly‑toned ideal.
  • Subtype 3: Muscle Dissatisfaction: Expressing perceived insufficient muscularity despite
  having a muscular physique.
Type 2: Relationship to Food:
  • Subtype 1: Rigid Food Rules: Obsessive macro/micronutrient tracking
  and restrictive dietary practices to rapidly gain muscle mass (bulking) or lose fat (cutting).
  • Subtype 2: Unsafe Food: Promotion of unconventional foods believed
  to enhance muscle growth.
  • Subtype 3: Cheat Meals: Large "reward" meals after restrictive dieting
  that reinforce binge–compensation cycles.
Type 3: Supplement Abuse:
  • Subtype 1: Anabolic Steroids: Normalization or
  endorsement of anabolic‑androgenic steroid use with downplayed risks.
  • Subtype 2: Legal APEDs: Overuse of legal supplements (e.g., creatine,
  protein, pre-workout) beyond recommended doses.
  • Subtype 3: Hormone Therapy: Downplaying risks and spreading misinformation
  about testosterone replacement therapy, promoted as a performance enhancer or "anti-aging" treatment
  without proper medical diagnosis.
Type 4: Relationship to Exercise:
  • Subtype 1: Excessive Exercise: Extreme exercise routines exceeding healthy limits despite
  injury or life interference.
  • Subtype 2: Predebting: Treating exercise as punishment or permission to eat.
  • Subtype 3: Maladaptive Coping: Using exercise as the sole coping mechanism to avoid
  emotional distress.
  • Subtype 4: Exercise‑Induced Functional Impairment: Prioritizing exercise
  over essential duties, harming daily functioning.
  • Subtype 5: Toxic Motivation: Demeaning communication that pressures
  unrealistic fitness standards via shaming or slurs.
Type 5: Relationship to Masculinity:
  • Subtype 1: Linking muscle‑building and exercise performance to male identity,
  sexuality, and self‑worth.
Type 6: Irrelevant:
  • Subtype 1: Content without muscle‑obsession, restrictive
  diets, supplement/AAS promotion, or extreme exercise. Includes dance trends,
  memes, travel/cooking vlogs, general wellness (e.g., yoga), and pure entertainment unrelated
  to body‑image narratives.
Here are some examples from the training data:
Example 1:
<video>
Caption: <Video Caption>
Audio transcript: <Audio Transcription from Whisper>
Embedded text: <OCR Text>
Classification: <Type>
Example 12:
<video>
Caption: <Video Caption>
Audio transcript: <Audio Transcription from Whisper>
Embedded text: <OCR Text>
Classification: <Type>
<video>
VIDEO_ID:
Caption:
Audio transcript:
Embedded text:
###TASK###
Classify the video into one of the following subtypes: Muscularity Self‑Objectification,
Leanness Self‑Objectification,
Muscle Dissatisfaction, Rigid Food Rules,
Unsafe Food, Cheat Meals,
Anabolic Steroids, Legal APEDs,
Hormone Therapy, Excessive Exercise,
Predebting, Maladaptive Coping,
Exercise‑Induced Functional Impairment,
Toxic Motivation,
Relationship to Masculinity, or Irrelevant.
Figure 9: Prompt used for few-shot inference for Task 2.
ID Batch Expertise Gender
A1 1 clinical psychology doctoral candidate specializing in male eating disorders and muscle dysmorphia intervention research. Male
A2 1 senior researcher in computational social science focused on social media dynamics and their implications for mental health. Female
A3 2 health communication professor using computational social network methods to study online health behaviors. Female
A4 2 medical student focusing on clinical and translational research in body image disorders and digital health interventions. Female
A5 3 clinical psychology doctoral candidate specializing in digital mental health and disordered eating prevention. Female
A6 3 communication doctoral candidate investigating how interactive media–driven social comparisons affect body image outcomes. Female
A7 4 clinical professor and psychiatrist studying the psychopathology and treatment of eating disorders and muscle dysmorphia. Male
A8 4 clinical professor and pediatric psychologist specializing in adolescent body image disorders. Female
A9 5 child and adolescent psychiatrist specializing in developmental psychopathology and body image disorders. Female
A10 5 clinical social worker specializing in child and adolescent body image disorders and psychosocial interventions in neuropsychiatric care. Female
A11 6 occupational therapist specializing in functional rehabilitation and psychosocial support for eating-disorder patients in a neuropsychiatric hospital setting. Female
A12 6 clinical dietitian specializing in nutritional assessment and dietary management for eating and body-image disorders. Female
A13 7 mental health nurse practitioner specializing in the assessment and treatment of child and adolescent psychiatric disorders. Female
A14 7 board-certified child and adolescent psychiatrist specializing in mood and behavioral disorders. Female
A15 8 clinical social worker with inpatient psychiatric experience in providing psychotherapy to adult patients. Female
A16 8 doctoral candidate specializing in detecting disordered behaviors in social media. Male
Table 11: Annotator assignments, areas of expertise, and gender.
Primary Secondary Definition Keywords
Relationship to Body Muscularity Self-Objectification Idealized imagery emphasizing muscular aesthetics as the primary source of value. shredded, swole, mensphysique, bodybuilding
Leanness Self-Objectification Self-evaluation against a lean, low-fat, highly-toned ideal promoted online. flat tummy, skinny men physique, small waist fitness
Muscle Dissatisfaction Expressing perceived insufficient muscularity despite fitness, fueling negative self-view. muscles never big enough, muscles not big enough, muscle dysmorphia, bdd men
Relationship to Food Rigid Food Rules Obsessive macro/micronutrient tracking and extreme bulking or cutting diets. aggressive cut, aggressive bulk, shredded diet, macro tracking
Unsafe Food Promotion of raw or unsafe foods believed to enhance muscle growth. liver king diet, raw meat diet for gym, dog food to gain muscles
Cheat Meals Large “reward” meals after restrictive dieting that reinforce binge–compensation cycles. cheatday food, cheatmeal
Relationship to Exercise Excessive Exercise Obsessive routines exceeding healthy limits despite injury or life interference. no rest day, david goggins mentality, train until failure, push your limit
Predebting Treating exercise as punishment or permission to eat, creating guilt cycles. exercise so I can eat, workout so I can eat, earn your food, train so I can eat
Maladaptive Coping Using exercise as the sole coping mechanism to avoid emotional distress. gym therapy, workout breakup, workout heartbreak, gym mental health, gym fixes everything
Exercise-Induced Functional Impairment Prioritizing exercise over essential duties, harming daily functioning. gym over everything, gym or nothing, gym over friends, skip school for gym
Toxic Motivation Demeaning communication that pressures unrealistic fitness standards via shaming or slurs. gym masculinity, aggressive gym motivation, they don’t know me son, get your ass to the gym
Supplements Anabolic Steroids Normalization or endorsement of anabolic-androgenic steroid use with downplayed risks. tren, anabolic stack, anabolic gear, steroids
Legal APEDs Overuse of legal supplements (e.g., creatine, protein, preworkout) beyond recommended doses. protein powder, whey, creatine, pre workout
Hormone Therapy Downplaying risks and spreading misinformation about testosterone replacement therapy. TRT for gains, testosterone, hormoneboost
Relationship to Masculinity Links muscle-building and exercise performance to male identity, sexuality, and self-worth. gym masculinity, be a man gym, embrace masculinity, alpha male gym
Irrelevant General fitness or lifestyle content unrelated to pro–bigorexia harm. fyp, tiktok, foryoupage, viral, funny, duet, trending, love, meme, followme, repost, new, awesome, music, cute, video, foryou, fun, diy, ootd, family, lifehack, photography, usa, college, travel, christmas, sport, party, popular, clip, movie, star, moment, tiktokviral, tiktokfamous, tiktokmusic, likeforfollow, recipe, quotes, TikTokChallenge, memories
Table 12: Primary and secondary categories of the Harm Taxonomy for Pro-Bigorexia Content on TikTok, with definitions and search keywords.
Video ID Subcategory Video Description Embedded Text Audio Transcript
7231300164307143979 Relationship to Body my lil frame me being a 5’8 male but with lean body and my little frame and my sweet little girl voice. It exudes something in people that…
7201534759347014958 Relationship to Masculinity #menshealth #menshealthtips #cynic hermes.the.cynic The reason that there is such a void of male role models for young men is because we literally do not have adult males in our population anymore. What do I mean by that? There’s no adult males. Like, there’s grown men everywhere, but they’re not actually physiologically adult human males…
7275865970306977067 Supplement Abuse this is totally healthy right? #recipes #creatine @days.of.j So I told a co-worker I was just taking my creatine in water, just water, because it was unflavored. And they told me you’re supposed to take it with like a meal or like sugar so then you can absorb it better…
7363842693211917611 Relationship to Food Liver King dinner! @liverking Pulled beef with hard shelled eggs Mmm It’s crazy. It actually tastes like the delicious pulled beef That’s in with the garlic oh and the onion and the lime juice booyah, yeah Sweet potatoes, the purple variety…
7342974339907603718 Relationship to Exercise I am girl MAKING THE GYM YOUR WHOLE PERSONALITY no train legs girls who make going to the gym their entire personality it is just so crazy to me I genuinely do not know how you guys do it I could never be one of those girls anyway someone I know just asked me what day it was and I replied with leg day
7498781444408429867 Irrelevant Two in #shorts #gymshorts @johnbshop0 ugly, ugly, ugly aura. Bro, if your shorts, especially if you’re hitting leg day, are below your knees, you’re genuinely cooked, bro. These, I got them here on the TikTok shop. This is a size medium. I’m five foot 10, 165 pounds, and these things fit so good…
Table 13: Example entries for Type classification showing captions and TikTok metadata. The actual videos can be found in the data attachment in the submission platform.
Video ID Subcategory Video Description Embedded Text Audio Transcript
7478843939676507423 Muscularity Self-objectification training hard or hardly training ? sounds of my workout on push day Yeah, we need some weight, should be a good one today. Bet you gotta hit me fast, you know old…
7231300164307143979 Leanness Self-objectification my lil frame me being a 5’8 male but with lean body and my little frame and my sweet little girl voice. It exudes something in people that…
7365667609435573509 Muscle Dissatisfaction #gymtok When u start working out to look better but now u have body dysmorphia fucking go let’s go I guess
7186715652642622762 Rigid Food Rules I gained 35lbs in 90 days #lifting #gym #bulking 35lbs in 90 days I’ve considered ever since 7th grade but change I gained 35 pounds in 90 days. I’ve considered myself a runner ever since 7th grade…
7332272427986111790 Unsafe Food Liver King Dinner to start the week liverking Mmm, steak and potatoes, sweet potato fritters, Liver King Chef Lionel, that’s how he calls them. So good. We got…
7398320286996663595 Cheat Meals NO MORE.. #gym #cheatday #gains us after cheat day turned into cheat month No more cheese fillets. No more McDonald’s. No more chicken wings. No more chicken snobby with some sour. No more…
7198309032384417070 Excessive Exercise Gym rats! #gym #gymtok #mothersoftiktok #thick #newyearsresolutions #muscles #gains #proteinshake #lawenforcement It’s time for us normal gym peeps to get our routines back!!! Listen, I got my ass the gym today and I didn’t want to go. So did you go today? Well…
Table 14: Example entries for Subtype classification showing captions and TikTok metadata (Part 1 of 2). The actual videos can be found in the data attachment in the submission platform.
Video ID Subcategory Video Description Embedded Text Audio Transcript
7372218805738589445 Predebting Exercise You CANNOT our train a bad diet #diet #fitness #exercise I CAN EAT WHATEVER I WANT AS LONG AS I EXERCISE YUP YOU HEARD ME WHILE EXERCISE IS IMPORTANT… I can eat whatever I want as long as I exercise. You said what? Yup, you heard me. While exercise…
7231695341077138730 Maladaptive Coping Gym Therapy #gymtherapy #gym #doordie #dobetter #military @shreddedvets Gym therapy is a term that was coined by Shredded Vets that describes the psychological effects of working out consistently…
7460046778524617991 Exercise-Induced Functional Impairment Bradman Best - I would skip school just to go to the gym. I was going to do whatever it took. I would skip school to guru the gym. lucky enough like when I was in year 9 and year 10 the Knights were giving me a day off…
7478474531875704094 Toxic Motivation put in the effort #gym #motivation #fitnessmotivation #workout TD PRESSURE It’s fucking different. Gotta really come kill me. I’m built different. I train different. I work different. I am different…
7313287497650326816 Anabolic Steroids #samsulek POV: Steroids gives you acne HAVE FUCKING I MAN LIKE PIMPLE SHIT COLLAB MIGHT PIMPLE UH ASIAN dude I think I have like a fucking I got like a cyst man you know like where it’s a…
7213153335786884394 Legal APEDs Sustenanceee #fyp #xyzbca #gymtok #creatine POV: you get caught using creatine Are there any drugs in this house? If there are, you better find them and give them to me immediately…
7190162324563496234 Hormone Therapy Day 5 on TRT #trt #lift #weightlifting #gym #workout Day 5 on TRT What up guys, it’s day five. So on day five, the doctor calls you with your full blood work done…
Table 15: Example entries for Subtype classification showing captions and TikTok metadata (Part 2 of 2). The actual videos can be found in the data attachment in the submission platform.